40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 24 00:42:07 2013, Jenne, Update, SUS, PRMI locked, ITMX pitch OpLev ringing up PRMI_locked_ITMXpitOpLevRingingUp.pdf
    Reply  Wed Apr 24 00:44:33 2013, Koji, Update, SUS, PRMI locked, ITMX pitch OpLev ringing up 
Message ID: 8482     Entry time: Wed Apr 24 00:44:33 2013     In reply to: 8481
Author: Koji 
Type: Update 
Category: SUS 
Subject: PRMI locked, ITMX pitch OpLev ringing up 

I tried to reproduce the locking situation described in this entry tonight.
The momentary lock was regularly seen but there was no stable lock.

I wonder why the actuators are always saturated. The feedback signals have the dominant component at ~400Hz.

It would also be nice if the servos have some immunity to gain fluctuation.

I didn't check how the situation of the AP table is. I'll look into some details tomorrow.

ELOG V3.1.3-